A skater spins at 2.50 revolutions per second. She slows with constant angular acceleration, stopping after 1.75 s. (a) What s her angular acceleration?


1
Expert's answer
2021-04-23T11:07:46-0400
"\alpha=\frac{\omega-\omega_0}{t}\\\alpha=2\pi\frac{0-2.5}{1.75}=-8.98\frac{rad}{s^2}"